OpenStudy (anonymous):

Nicole Is Taking Some Friends On A Picnic. They'll Need To Follow A Path To Get To Picnic Spot. A Map Of The Path Is Based On A Scale Of 1:10,000, In CM. If The Path Is 20 CM On The Map, How Many Kilometers Is The Actual Path ?

OpenStudy (anonymous):

@AbhimanyuPudi

OpenStudy (anonymous):

20 * 10,000 = 200,000 cm km = 100000 cm 200,000/100,000 = 2km

OpenStudy (anonymous):

1cm path is 10,000cm in graph 20cm path is 20 x 10,000 = 2,00,000cm 1km = 100000cm So, 1cm = (1/100000)km 2,00,000cm = (200000/100000)km = 2km

OpenStudy (anonymous):

Okay So Is This Correct For My Next Question 2.1*25000=52500 CM KM=100000 CM 52500/100000=0.525

OpenStudy (anonymous):

If the scale is 1:25,000 in cm and the path is 2.1cm, then you are right!!

OpenStudy (anonymous):

Thanks Thats It

OpenStudy (anonymous):

welcome:)